OpenStudy (anonymous):

How can I integrate this?\[\int_{0}^{2}\int_{0}^{1}(2x+y)^8dxdy\]

OpenStudy (jamesj):

Just integrate it calmly as you might expect. The inner integral, that wrt x, is just \[ \left.\frac{1}{18} (2x + y) \right|_{x=0}^1 \] Evaluate that then integrate wrt y.

OpenStudy (jamesj):

1/18 . (2x + y)^9

OpenStudy (anonymous):

oh i see it now. many thanks JamesJ